Q1. Suppose Random variable X is # of jobs students applied with minimum 0 to maximum 23 that are uniformly distributed.

  1. Find P(2<x<18)
Answer #1

Solution :

Given that,

a = 0

b = 23

P(c < x < d) = (d - c) / (b - a)

P(2 < x < 18) = (18 - 2) / (23 - 0)

= 16 / 23 = 0.6957

P(2 < x < 18) = 0.6957